The following water quality data are available for a community

Water quality

Ca2+ = 60 mg/l (as species)

Mg2+ = 40 mg/l (as species)

HCO3- = 220 mg/l as CaCO3

CO2 = 7 mg/l (as species)

pH = 7.6

Final desired hardness = 80 mg/l as CaCO3.

The flow rate of the treatment plant is 0.24 m3/s.

a. Calculate the total, carbonate, and non-carbonate hardness

b. Calculate the amount of lime (as mg/l of CaO) needed to soften the water using the standard lime-soda treatment process. Assume that the lime is 90% pure.

c. Calculate the amount of soda ash needed (as mg/l of Na2CO3) to soften the water using the standard lime-soda treatment process. Assume that the soda ash is 94% pure.
Answer #1

SOLUTION:

Hardness (in mg/L as CaCO3) = [mass of Ca2+ (in mg/L) × 50]/ (E.Wt. of Ca2+)

Here, E.Wt. of Ca2+ = (40g/mole)/2=20 g/mole

E.Wt. of Mg2+ = (24.4g/mole)/2=12.2 g/mole

E.Wt. of CO2 = (44g/mole)/2=22 g/mole

a.

Compound      mg/L as spices EW/EW      mg/L as CaCO3

               Ca2+ 60.0 20.0 150.0

               Mg2+ 40.0 12.2 163.93

               HCO3                  ---                    ‑‑-                   220.0

CO2 7.0 22 15.91

TH = 150+163.93+15.91 = 329.84 mg/L as CaCO3

CH = HCO3 = 220.0 mg/L as CaCO3

NCH = TH ‑ CH = 329.84 ‑ 220.0 = 109.84 mg/L as CaCO3

b. CO2+ = 7 mg/l x 1.27 = 8.89 mg/l as CaO

HCO3= 220 mg/l x 0.56 = 123.2 mg/l as CaO

Mg2+ =40 mg/l x 2.31 = 92.4 mg/l as CaO

  Ca2+ = 60 mg/l x 0.56 = 33.6 mg/l as CaO

TOTAL = 258.09 mg/l as CaO

Actual Lime Dose = 258.09 mg/l / 0.9 = 286.77 mg/l

c. Non carbonate Hardness = 109.84 mg/l as Caco3

Soda Ash Dosage = 109.84 mg/l x 1.06 = 116.43 mg/l as Na2CO3

Actual Dosage = 116.43 mg/l / 0.94 = 123.86 mg/l as Na2CO3